OpenStudy (anonymous):

Simplify quantity 4 x squared plus 12 x minus 16 all over quantity 2x plus 10 over quantity 6 x plus 24 over quantity x squared plus 9x plus 20 Would someone please explain this to me??

OpenStudy (campbell_st):

well it looks like this \[\frac{\frac{4x^2 + 12x + 16}{2x +10}}{\frac{6x + 24}{x^2 + 9x + 20}}\] the rule for dividing by a fraction is flip the denominator and multiply so you have \[\frac{4x^2 + 12x + 16}{2x + 10} \times \frac{x^2 + 9x + 20}{6x + 24}\] to get an answer, you'll need to factor the expressions. then cancel common factors and lastly multiply whats left. Hope this helps

OpenStudy (campbell_st):

oops it should be -16 not plus... \[\frac{4x^2 + 12x - 16}{2x + 10} \times \frac{x^2 + 9x + 20}{6x + 24}\]

OpenStudy (anonymous):

How can I factor 9 and 2 if they have no common factors?

OpenStudy (campbell_st):

well start this way \[\frac{4(x +4)(x -1)}{2(x + 5)} \times \frac{(x + 4)(x + 5)}{6(x + 4)}\]

OpenStudy (anonymous):

Okay I got it. Thanks!!!

OpenStudy (anonymous):

was the answer 3(x-1)/x+4 ?